Parallel Adaptive Multigrid Methods in Linear and Nonlinear Structural Mechanics

Knut Eckstein

Graduiertenkolleg Kontinua und Strömungen, ISD, Universität Stuttgart

This talk presents the application of parallel adaptive multigrid methods to linear and nonlinear problems in structural mechanics. The multigrid method is employed as an iterative solver on unstructured locally refined 2D grids discretizing plane elasticity and plasticity problems with triangular and quadrilateral elements.

Subtopics include the discussion of coarse grid assembly techniques in the nonlinear case, the Newton-Raphson linearization, the isotropic and kinematic hardening laws modelled in the nonlinear material formulation as well as indicators used in adaptive grid refinement and dynamic load balancing schemes.